### v1

# Arena: A Patch-of-Interest ViT Inference Acceleration System for Edge-Assisted Video Analytics

Abstract.
The advent of edge computing has made real-time intelligent video analytics feasible. Previous works, based on traditional model architecture (e.g., CNN, RNN, etc.), employ various strategies to filter out non-region-of-interest content to minimize bandwidth and computation consumption but show inferior performance in adverse environments. Recently, visual foundation models based on transformers have shown great performance in adverse environments due to their amazing generalization capability. However, they require a large amount of computation power, which limits their applications in real-time intelligent video analytics. In this paper, we find visual foundation models like Vision Transformer (ViT) also have a dedicated acceleration mechanism for video analytics. To this end, we introduce Arena, an end-to-end edge-assisted video inference acceleration system based on ViT. We leverage the capability of ViT that can be accelerated through token pruning by only offloading and feeding Patches-of-Interest (PoIs) to the downstream models. Additionally, we employ probability-based patch sampling, which provides a simple but efficient mechanism for determining PoIs where the probable locations of objects are in subsequent frames. Through extensive evaluations on public datasets, our findings reveal that Arena can boost inference speeds by up to 1.58$\times$ and 1.82$\times$ on average while consuming only 54% and 34% of the bandwidth, respectively, all with high inference accuracy.
Video Analytics System, Edge Computing, Vision Transformer
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Computer vision tasksccs: Computer systems organization Real-time system architecture
1. Introduction
Today, IoT cameras are generating an unprecedented volume of video data daily. According to predictions by Hailo (Hailo, 2024), the global market is anticipated to see approximately 180 million shipments of embedded video surveillance devices by 2025. These cameras power a plethora of video analytics applications such as traffic management (Tong et al., 2021; Li et al., 2023), autonomous driving (Shi et al., 2022; Zhang et al., 2021b), security surveillance (Yang et al., 2023; Yi et al., 2020), etc.
The massive data generated by IoT cameras far exceeds their limited computing capabilities. A straightforward way is to offload the captured data to edge servers. These servers execute intensive downstream inference tasks, usually based on deep neural networks, and deliver useful visual feedback to users. Previous works have explored the spatial and temporal redundancy of video content to design systems that reduce the bandwidth usage between cameras and edge servers. Efforts are made either to extract and transmit regions of interest (RoIs) from video frames (Liu et al., 2022b; Zhang et al., 2021a) or to design a module that filters out similar frames (Tchaye-Kondi et al., 2022; Chen et al., 2015). Though these methods have optimized the transmission and computation, their downstream models based on Convolutional Neural Network (CNN) show inferior performance in the live video data, because of the poor generalization ability of CNN.
Figure 1. Arena: our patch-of-interest ViT inference acceleration system for edge-assisted video analytics. Due to the limited computing power of the camera, the extracted patches-of-interest are offloaded to an edge server for processing with its more powerful GPUs. MTPs stands for Memory Token Pools.
[…]
Our Edge-assisted computation offloading in video analytics.
In recent years, visual foundation models based on transformer architecture (Dosovitskiy et al., 2020; Liu et al., 2022a, 2021; Li et al., 2022), exemplified by Vision Transformer (ViT) (Dosovitskiy et al., 2020), have surpassed CNN-based methods (Girshick, 2015; Redmon and Farhadi, 2018; Cai and Vasconcelos, 2019; Tian et al., 2019) in most downstream tasks, including image classification (Deng et al., 2009), object detection (Lin et al., 2014) and object tracking (Milan et al., 2016). The strengths of the ViT lie in its exceptional feature extraction capabilities and generalization. Through its multi-head self-attention mechanism (MSA) (Vaswani et al., 2017), ViT can more effectively capture global information within the input, achieving superior performance in various aspects. Furthermore, the ”pre-training + fine-tuning” paradigm has also standardized the implementation process of video analytics, enabling models to be rapidly deployed in entirely new specific scenes using only a few samples. However, the computational demands of its MSA increase quadratically with the number of patches. This intense computational requirement, especially in high-resolution video (e.g., 1080p) analytics scenarios, impedes the practical application deployment of ViT-based models. Although there exits works on efficient ViT that utilize token pruning (Rao et al., 2021; Liang et al., 2022), merging (Bolya et al., 2022; Zeng et al., 2022), and fusion (Kim et al., 2024) techniques to reduce computation, these approaches still require a full image as input and fail to adapt to real-time intelligent video analytics.
To fill this gap, we propose Arena, a Patch-of-Interest (PoI) ViT inference acceleration system for edge-assisted video analytics. We find that ViT also has a dedicated acceleration mechanism for video analytics. As shown in Figure 1, first, the camera transmits a full video frame (referred to as a keyframe) to the edge server, followed by conducting inference on it, where we establish and maintain memory token pools. Subsequently, we utilize a Probability-based Patch Sampling (PPS) mechanism to determine the PoIs for the next frame (referred to as a Non-keyframe). Therefore, only PoIs, usually occupying a small portion of the frame, are required to transmit to the edge server. Then, a model based on the ViT backbone performs inference merely using these PoIs. Finally, a Memory Feature Reconstruction (MFR) module is employed to restore the complete feature maps, making it available for a dense prediction task head. Arena enhances overall system throughput and computational efficiency twofold. On the one hand, transmitting only partial patches significantly reduces bandwidth consumption and transmission latency. On the other hand, feeding PoIs to the ViT backbone and reconstructing features from memory tokens reduce computational overhead and inference time while almost maintaining the same accuracy.
We develop and deploy a prototype on a testbed that uses an NVIDIA Jetson as a camera device and an Ubuntu desktop as an edge server running real video analytics workloads. Our results show that Arena can accelerate the inference up to 1.58×1.58$\times$1.58 × and 1.82×1.82$\times$1.82 × on average by only using 54% and 34% bandwidth while keeping the inference accuracy sacrifice within an acceptance range.
The remainder of this paper is organized as follows. We present the preliminaries and motivation in Section 2. The design of Arena is introduced in Section 3, followed by the system implementation in Section 4. We conduct extensive experiments in Section 5, and the related work is reviewed in Section 6. Finally, we conclude this paper in Section 7.
2. Preliminaries And Motivation
We first introduce the preliminaries of Vision Transformer (Sec. 2.1). Next, we reveal that video data contains a significant amount of redundancy. Substantial bandwidth can be saved by only transmitting the PoIs (Sec. 2.2). ViT is well-suited for unstructured data because it can process variable-length sequences as inputs. This capability allows for filtering out data at the beginning of the pipeline and accelerates the computation (Sec. 2.3), which gives us insights on how to accelerate ViT-based video analytics systems.
2.1. Preliminaries of Vision Transformer
Given an input image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$, ViT splits it into a sequence of flattened 2D patches $\mathbf{x}_{p}=\{x_{p,1},\cdots,x_{p,N}|x_{p,i}\in\mathbb{R}^{P^{2}C}\}$. Here, $H\times W$ represents the resolution of the frame, $C$ is the number of channels, $P$ denotes the size of each patch, and $N=(H\times W)/P^{2}$ is the number of patches after splitting. After that, ViT constructs initial tokens as
(1) $\tilde{z}_{0}=\left[x_{class},x_{p,1}\mathbf{E};x_{p,2}\mathbf{E};\ldots;x_{p,N}\mathbf{E}\right],$
where $x_{class}$ is a learnable embedding that prepends into flattened sequences and $\mathbf{E}\in\mathbb{R}^{P^{2}C\times D}$ represents the parameters of the linear projection, and $D$ denotes the embedding dimension of each token. Moreover, ViT adds position embeddings $\mathbf{E}_{pos}$ on $\tilde{z}_{0}$ to retain its positional awareness as
(2) $z_{0}=\tilde{z}_{0}+\mathbf{E}_{pos},\qquad\mathbf{E}_{pos}\in\mathbb{R}^{(N+1)\times D}.$
In the end, the transformer encoder (Vaswani et al., 2017) consisting of MSA, MLP blocks, and LayerNorm (LN) is adopted to update $z_{0}$ with $L$ iteration and output the final image representation $\mathbf{y}$. The formulations are
(3) $z_{\ell}^{\prime}$ $=\operatorname{MSA}\left(\operatorname{LN}(z_{\ell-1})\right)+z_{\ell-1}^{\prime},$ $\ell=1\ldots L,$
$z_{\ell}$ $=\operatorname{MLP}\left(\operatorname{LN}(z^{\prime}{}_{\ell})\right)+z_{\ell}^{\prime},$ $\ell=1\ldots L,$
(4) $\mathbf{y}=\mathrm{LN}(\mathbf{z}_{L}^{0}).$
2.2. Redundancy in Video Inference Data
In edge-assisted video analytics, high-resolution videos often include much redundant data. Particularly, the video frame is divided into multiple patches of certain pixels (e.g., $16\times 16$), among which those containing the objects are referred to as PoIs. Conversely, the rest of the frame is dominated by the background (e.g., buildings and sky) and other irrelevant objects (Jiang et al., 2021). Table 1 shows the redundancy of the first five cameras from two popular real-world video datasets. It is evident that in most videos, the proportion of PoI does not exceed 25%. Additionally, in some specific scenes, PoIs constitute no more than 5% of the whole frame. By only transmitting the PoIs from the camera to the edge server, most videos can achieve a bandwidth usage reduction exceeding 80%, with potential savings reaching up to 92.01%. This discovery motivates us to seek a more fine-grained method for extracting PoIs that allows transmitting as few patches as possible while still meeting the accuracy requirements.
Table 1. Redundancy in video inference data on MOT17Det (Milan et al., 2016) and AIC22 (Naphade et al., 2022) datasets.
Scene Name (# Frame) # Object PoI Prop. Latency Bandwidth
MOT17-02 (600) 9393 17.10% 90.11% 67.29%
MOT17-04 (1050) 39821 22.59% 86.55% 62.60%
MOT17-05 (837) 4767 69.21% 48.01% 21.70%
MOT17-09 (525) 3745 24.46% 85.67% 57.72%
MOT17-10 (654) 10560 13.76% 89.14% 56.64%
AIC22-c001 (1955) 3310 3.79% 86.76% 92.01%
AIC22-c002 (2110) 6322 4.77% 85.24% 90.86%
AIC22-c003 (1996) 6677 8.69% 85.40% 87.14%
AIC22-c004 (2110) 6015 4.88% 84.92% 88.05%
AIC22-c005 (2110) 12199 5.40% 74.60% 88.75%
  • # represents “The number of”;
  • $\triangle$ The average proportion of the total PoIs in each frame;
  • $\diamondsuit$ Average proportion of inference latency saved by only computing PoIs on ViT backbone;
  • $\star$ Saved bandwidth proportion if only transmits the PoIs.
2.3. Accelerating ViT via Patch Prunning
Figure 2. Pruning patches can accelerate ViT backbone inference. We evaluate the impact of pruning 25%, 50%, and 75% tokens on (a) inference latency and (b) GFlops using ViT-base across videos of 1080p, 720p, and 480p resolutions. Similar trends were found in ViT-Base.
One of the desirable advantages of ViT is its ability to allow a flexible length of patch input. Specifically, the total computational complexity of one MSA module and one MLP is $\mathcal{O}\left(12ND^{2}+2N^{2}D\right)$. Obviously, reducing the number of patches can linearly or even quadratically reduce the operations. Figure 2 illustrates the (a) average inference latency and (b) average GFlops per frame when different proportions of patches are pruned across videos of various resolutions. It is noteworthy that using ViT-small on 1080p videos and pruning 75% patches resulted in a reduction of latency by 62ms, only 11.76% of the original latency. Additionally, the GFlops are linearly related to the number of input patches, showing a reduction of approximately 74.97% if pruning 75% patches. Furthermore, we confirm this on video datasets in Table 1. The Latency column reports the average inference latency saved by only feeding the PoIs to the ViT backbone for each frame. Most videos can achieve over an 80% acceleration, with the highest reaching up to 90.11%. This observation shows the potential opportunities for accelerating ViT inference by pruning the number of input patches.
However, naively pruning patches is not feasible for dense prediction tasks, such as object detection, since they typically require full-sized feature maps. How to adapt to downstream models while enjoying the inference acceleration brought by patch pruning presents a significant challenge to the design of Arena.
3. Arena Design
Figure 3. Given $K$ continuous frames $\{\hat{\mathbf{x}}^{1},\mathbf{x}^{2},\ldots,\mathbf{x}^{K}\}$ in an interval, Arena periodically operates in two distinct phases: keyframe inference (Left) for the first frame $\hat{\mathbf{x}}^{1}$ and non-keyframe inference (Right) for the rest of the frames. Both two phases use the same network architecture with shared weights. Notably, we split the frame into nine patches only for demonstration.
3.1. Overview
Arena is an edge-assisted, real-time inference acceleration system tailored for ViT-based models. It aims to reduce communication and computation overheads by pruning the number of patches offloaded to the server while maintaining high accuracy. Arena comprises a camera and an edge server. The camera, equipped with limited computational power, continuously captures video frames, and the edge server deploys a ViT-based detector to perform inference. In every $K$ frames, the first frame is designated as the keyframe, and the subsequent $K-1$ frames are considered as non-keyframes. Arena operates in two distinct phases: keyframe inference (Sec. 3.2) and non-keyframe inference (Sec. 3.3), then alternating periodically between them as illustrated in Figure 3. In keyframe inference, Arena first performs full frame inference by detector and caches intermediate tokens into two memory token pools. In the non-keyframe inference phase, it queries memory tokens from the pools and reconstructs complete image-wise features to accommodate unstructured samples with variable-length patch sequences. Moreover, we introduce PPS to identify the PoIs in non-keyframes and provide feedback to the camera (Sec. 3.4). Finally, Sec. 3.5 describes our training approach in a video-training manner. Notably, both phases employ the same network and shared weights. We take a sequence of frames $\{\hat{\mathbf{x}}^{1},\mathbf{x}^{2},\ldots,\mathbf{x}^{K}\}$ as an example, where $\hat{\mathbf{x}}^{1}$ denotes the keyframe and $\mathbf{x}^{2},\ldots,\mathbf{x}^{K}$ as non-keyframes, and detail our above-mentioned mechanisms in the following paragraphs.
3.2. Keyframe Inference
In the keyframe inference phase, the camera transmits a full frame to the edge server. For keyframe $\hat{\mathbf{x}}^{1}$, the ViT detector splits it into a sequence of non-overlapping patches with resolution $P\times P$, denoted as $\hat{\mathbf{x}}_{p}^{1}=\{\hat{x}^{1}_{p,1},\cdots,\hat{x}^{1}_{p,N}|\hat{x}^{1}_{p,i}\in\mathbb{R}^{P^{2}C}\}$. After that, following the procedure of ViT, Arena constructs the initial tokens $\hat{\tilde{z}}_{0}^{1}$ as
(5) $\mathcal{M}_{a}\leftarrow\hat{\tilde{z}}_{0}^{1}=[\hat{x}^{1}_{p,1}\mathbf{E};\hat{x}^{1}_{p,2}\mathbf{E};\ldots;\hat{x}^{1}_{p,N}\mathbf{E}],$
where the initial tokens $\hat{\tilde{z}}_{0}^{1}$ without positional encoding will be cached in the memory token pool $\mathcal{M}_{a}$. After that, position embeddings $\mathbf{E}_{pos}$ are added to the initial tokens to retain positional information as
(6) $\hat{z}_{0}^{1}=\hat{\tilde{z}}_{0}^{1}+\mathbf{E}_{pos},\qquad\mathbf{E}_{pos}\in\mathbb{R}^{N\times D}.$
Then, all tokens are processed through the $L$ layers transformer encoder to get the final output $\hat{z}_{L}^{1}$, which will be cached in the second memory token pool $\mathcal{M}_{b}$. The formation is:
(7) $\hat{z}_{\ell}^{1\prime}$ $=\operatorname{MSA}\left(\operatorname{LN}(\hat{z}^{1}_{\ell-1})\right)+\hat{z}^{1}_{\ell-1},$ $\ell=1\ldots L,$
$\hat{z}_{\ell}$ $=\operatorname{MLP}\left(\operatorname{LN}(\hat{z}_{\ell}^{1\prime})\right)+\hat{z}_{\ell}^{1\prime},$ $\ell=1\ldots L,$
(8) $\mathcal{M}_{b}\leftarrow\hat{z}_{L}.$
Notably, memory token pools $\mathcal{M}_{a}$ and $\mathcal{M}_{b}$ will be updated in every frame inference and re-initialized at the next keyframe inference phase. Since the tokens of the keyframe are complete, we directly feed $\hat{z}_{L}$ into the MFR module. MFR is a single-layer transformer decoder (Vaswani et al., 2017) to build the final tokens and reshape to the strongest feature map $f_{3}$.
We build a multi-scale feature pyramid $\{f_{1},f_{2},f_{3},f_{4}\}$ directly from intermediate tokens. $f_{3}$ and $f_{4}$ contain low-level information, while $f_{1}$ and $f_{2}$, after processing through the transformer encoder, contain high-level semantic information. Therefore, we combine these two sets of feature maps to enhance the effectiveness of object detection. Specifically, we use the token sequence before encoder (i.e., $\hat{\tilde{z}}_{0}^{1}$) and after MFR (i.e., $f_{3}$) as features, which are both at $1/16$ scale relative to the original frame. For $\hat{\tilde{z}}_{0}^{1}$, we use two sets of de-convolution operations to obtain feature maps $f_{1}$, $f_{2}$ at $1/4$ and $1/8$ scale. For $f_{3}$, we use the convolution of stride 2 to generate a feature map $f_{4}$ at $1/32$ scale. The formulations are
(9) $f_{1}=\text{DeConV}_{1}(\hat{\tilde{z}}_{0}^{1}),\qquad f_{2}=\text{DeConV}_{2}(\hat{\tilde{z}}_{0}^{1}),$
(10) $f_{3}=\text{MRF}(\hat{z}^{1}_{L}),\qquad f_{4}=\text{ConV}(f_{3}).$
Finally, $\{f_{1},f_{2},f_{3},f_{4}\}$ from the ViT backbone form a multi-scale feature pyramid, which is then fed into two-stage detectors, such as Faster R-CNN (Ren et al., 2015) or Cascade R-CNN (Cai and Vasconcelos, 2019), to generate the bounding boxes and categories of the objects.
3.3. Non-keyframe Inference with Memory Feature Reconstruction
During non-keyframe inference, the PPS first identifies $N^{\prime}$ PoIs using the last frame detection results (e.g., results from $\hat{\mathbf{x}}^{1})$. Then, the camera only extracts the PoIs out of the frame and offloads them to the edge server, which is denoted as $\mathbf{x}_{p}^{2}=\{x^{2}_{p,1},\ldots,x^{2}_{p,N^{\prime}}|x^{2}_{p,i}\in\mathbb{R}^{P^{2}C}\}$. Here $N^{\prime}\ll N$ due to the fact that objects only occupy a very small portion of the frame. After linear projection for $\mathbf{x}_{p}^{2}$, we get a sparse token sequence $\tilde{z}_{0,sp}^{2}$. Subsequently, we add position embeddings corresponding to the original frame position to each token as
(11) $z_{0,sp}^{2}=\tilde{z}_{0,sp}^{2}+\mathbf{E}^{\prime}_{pos,sp},\qquad\mathbf{E}^{\prime}_{pos,sp}\in\mathbb{R}^{N^{\prime}\times D}.$
Next, $z_{0,sp}^{2}$ is fed into the transformer encoder to compute $z_{L,sp}^{2}$ following Eq. (7).
To restore the complete token sequence from the sparse one, we need to reuse the tokens in memory token pools. We extract tokens from non-PoI regions in $\mathcal{M}_{a}$ and combine them with the PoIs token $\tilde{z}_{0,sp}^{2}$ to restore a complete token sequence $\tilde{z}_{0}^{2}$. Similarly, we extract tokens from non-PoI regions in $\mathcal{M}_{b}$ and combine them with $z_{L,sp}^{2}$ to restore $z_{L}^{2}$. For example, as illustrated in Figure 3, the edge server only receives and processes tokens at positions 1, 2, 3, 5, and 6. Then, we insert the tokens from the remaining positions of $\mathcal{M}_{a}$ and $\mathcal{M}_{b}$ (i.e., 4, 7, 8, and 9) into the processed tokens $\tilde{z}_{0,sp}^{2}$ and $z_{L,sp}^{2}$ to restore complete sequences, respectively. Meanwhile, we introduce MFR to enhance the PoI tokens with background information from historical frames with minimal computational overhead. Finally, we follow the Eq. (9) and Eq. (10) to get the multi-scale feature pyramid $\{f_{1},f_{2},f_{3},f_{4}\}$.
In the end of inference phase, memory token pools $\mathcal{M}_{a}$ and $\mathcal{M}_{b}$ are updated by
(12) $\mathcal{M}_{a}\leftarrow\tilde{z}_{0}^{2},\mathcal{M}_{b}\leftarrow z_{L}^{2}.$
3.4. Probability-based Patch Sampling
Figure 4. Our probability-based patch sampling mechanism. The blue patch indicates that it is highly likely to be determined as an PoI. Zoom in for the best view.
How to determine the PoIs in the next frame is a critical challenge. Too few patches may cause the model to miss detections, while too many patches could transmit unnecessary background information and add extra computational overhead. To address this issue, we design a lightweight and effective probability-based patch sampling mechanism that leverages history information to determine the PoIs. Our insight is that the region where the object appears in the next frame will be near the bounding box of the previous frame with high probability.
As shown in Figure 4, the two frames are adjacent. In the first frame $\mathbf{x}^{k}$, we have obtained the bounding box of the objects (box in red), and all patches occupied by this bounding box (patches in red) are considered as the PoIs of the previous frame. Next, we expand the PoIs region by appending $m$ (referred to as the expanded bbox margin) rows and columns around the perimeter, including the top, bottom, left, and right edges (patches in yellow). These patches are regarded as the sampling region $R_{s}$ for the current frame $\mathbf{x}^{k+1}$ (patches in blue). We use inter-frame differences to determine the regions that need additional transmission to alleviate the miss of new objects that enter the scene for the first time. Specifically, we convert $\mathbf{x}^{k}$ and $\mathbf{x}^{k+1}$ to greyscale images and compute the pixel difference $f_{d}$ between the two greyscale images. Next, we divide the greyscale image $f_{d}$ into non-overlapping $16\times 16$ patches, identifying patches outside the $R_{s}$ with pixel differences exceeding the threshold $F$. The new objects are likely to be found in these patches, which are also marked as sampling regions. Finally, we randomly sample $p\%$ of patches from the sampling region as PoIs because our backbone is not sensitive to whether the object area is complete, and the background can be reconstructed using MFR.
3.5. Training Procedure
We fine-tune the whole detector based on a pre-trained ViT backbone in a video-training manner. During fine-tuning, for each training batch, we input a pair of adjacent frames $\{\hat{\mathbf{x}}^{k},\mathbf{x}^{k+1}\}$ to the model to simulate our keyframe and non-keyframe inference procedure. Both frames use identical preprocess pipeline, including data augmentation techniques such as cropping and flipping. After that, we take the first frame $\hat{\mathbf{x}}^{k}$ as the keyframe and feed all patches from $\hat{\mathbf{x}}^{k}$ into the ViT backbone to initialize the memory token pools $\mathcal{M}_{a}$ and $\mathcal{M}_{b}$. Since directly using predicted bounding boxes as results may disturb the proper detection in the next frame during training, we use the ground truth bounding boxes instead of predicted ones from the first frame as the detection results. Finally, for the second frame $\mathbf{x}^{k+1}$, we follow Sec. 3.4 to get the PoIs and Sec. 3.3 to conduct a non-keyframe inference. It’s worth noting that we only use the inference results of the second frame for back-propagation and network updates. In addition, we utilize the pre-trained weights of ViT backbone derived from masked image modeling techniques (He et al., 2022; Ren et al., 2023).
4. Implementation
We develop a prototype of Arena on an edge server equipped with two Intel(R) Xeon(R) Silver 4310 CPUs, 128 GB of RAM, and two NVIDIA GeForce RTX 4090 GPUs (16384 CUDA Cores) , each with 24 GB of VRAM. An NVIDIA Jetson AGX Orin running Ubuntu 22.04 LTS serves as the camera. The edge server and camera are connected with the TP-LINK TL-WDR5620 through a 2.4GHz WiFi network. The bandwidth is 93.9 Mbps, measured by iperf. Arena performs preprocessing on the camera to obtain the PoIs that need to be transmitted. The camera utilizes the HTTP protocol and creates request sessions to send POST requests to the server. The server returns the position information for PoIs of the next frame to the camera in JSON format. The ViT detector is implemented with Python and trained using MMTracking (Contributors, 2020) framework on the edge server. The PPS is implemented with C++ on the camera. Figure 5 shows our experimental platform.
Figure 5. Snapshot of our experimental evaluation hardware platform, including an edge server with GPUs and an NVIDIA Jetson AGX as the camera.
5. Evaluation
In this section, we conduct comprehensive experiments to evaluate the performance of the proposed Arena system. We start with the experimental setup and then compare Arena to baseline methods on three metrics. Next, we give insights into the accuracy and bandwidth trade-off of Arena. Finally, we report the model analysis.
5.1. Experimental Setup
5.1.1. Performance Metrics
The performance metrics we evaluate include accuracy, bandwidth usage, and end-to-end latency.
$\bullet$ Accuracy: In the object detection task, we evaluate detection accuracy using mAP@0.5, where an object is considered detected if the intersection over union exceeds 0.5. This metric is commonly employed in video analytics research.
$\bullet$ Bandwidth Usage: Arena aims to reduce the size of the offloaded data. To verify this, we use the data size of the frame sent to the edge server to measure the bandwidth usage. Since the bandwidth consumption is different between datasets, we report values after normalization.
$\bullet$ End-to-End Latency: Arena can accelerate the whole video analytics system in multiple aspects, so we compare the average end-to-end latency. End-to-end latency is defined as the time from the capture of a frame to the completion of inference, including the latency for frame preprocessing on the camera, transmission delay, and inference time.
5.1.2. Datasets
We evaluate Arena using the following two datasets.
$\bullet$ MOT17-Det (Milan et al., 2016): It contains moving or stationary street views with pedestrian bounding box labels from indoor and outdoor settings across 11 distinct regions. The resolutions of these videos are either $640\times 480$ or $1920\times 1080$. We split the initial 75% of each video sequence to constitute our training set, with the remaining 25% serving as the validation set.
$\bullet$ AI City Challenge 2022 (AIC22) (Naphade et al., 2022): AIC22 is a real-world multi-camera traffic dataset including 42,683 frames and 88,820 vehicle bounding box labels. It contains video footage from a total of 59 surveillance cameras situated at five traffic intersections. We select videos from three intersections to constitute our training set while setting the videos from the remaining intersections as the validation set.
5.1.3. Arena’s Configurations
We employ Faster R-CNN (Ren et al., 2015) as the two-stage detector framework and ViT-small as the backbone, which features a 12-layer transformer encoder and $D=384$. MFR is a one-layer transformer decoder that comprises merely 2.07M (5% of total) parameters and requires approximately 17.16 (1.9% of total) GFlops. The backbone pre-trained weights are from (Ren et al., 2023). Unless specifically stated, the keyframe interval $K=5$, $p=0.9$, and $F=200$. For the expanded bbox margin $m$, we set it to $1$ for the MOT17Det dataset and $3$ for the AIC22 dataset.
5.1.4. Baselines
We compare Arena with the following baselines.
$\bullet$ Full-frame Detector (FD): The camera consistently transmits the full frame to the edge server, which employs a single-frame ViT detector for inference without additional training techniques.
$\bullet$ Faster R-CNN Detector (Ren et al., 2015) (FRD): Similar to FD, but the edge server employs a Faster R-CNN detector with ResNet-50.
$\bullet$ DDS (Du et al., 2020): DDS transmits low-quality videos to the edge server to identify RoIs, which are fed back to the camera. The camera subsequently transmits these regions in high quality for second-round inference. We simulate different qualities by adjusting the resolution.
$\bullet$ Reducto (Li et al., 2020): Reducto filters out frames with minimal impact on the accuracy by comparing low-level video feature differences between two adjacent frames at the camera to reduce bandwidth consumption. We employ Edge feature and utilize training sets to profile the optimal threshold.
$\bullet$ ELF (Zhang et al., 2021a): ELF uses attention-based LSTM to predict the bounding box shifts frame by frame. It then employs a frame partitioning method to offload the region proposals to multiple edge servers. We utilize the official implementation and configure it for three servers with parallel offloading and inference.
5.2. Overall Performance of Arena
Figure 6. The normalized bandwidth usage of different methods on two datasets.
Figure 7. The accuracy of different methods on two datasets. Arena can maintain accuracy losses within 1% and 4%.
Figure 8. The average end-to-end latency per frame of different methods on two datasets. End-to-end latency includes a breakdown of preprocessing, transmission, and inference time.
In Figure 6 and Figure 7, we report the normalized bandwidth usage and accuracy of different methods on two datasets. Figure 8 displays the end-to-end latency of each method, including a breakdown of preprocessing, transmission, and inference time. Among these baselines, Arena achieves the best trade-off among bandwidth, latency, and accuracy in both datasets. FD and FRD always transmit the full frame for inference, representing the upper bound for bandwidth consumption. They achieve the highest accuracy on their models but also result in the longest end-to-end latency. DDS achieves high bandwidth efficiency by transmitting only a few regions in high quality but inevitably incurs additional delay due to two rounds of inference. Reducto struggles to adapt to object detection tasks, as removing too many frames significantly reduces its accuracy. ELF maintains low bandwidth consumption and end-to-end latency, but this is contingent upon the premise of parallel offloading and inference by three servers. In contrast, Arena substantially saves bandwidth usage by pruning patches in non-keyframes, thus reducing both transmission delay and the model inference time. Meanwhile, Arena utilizes historical frame information to reconstruct features, minimizing the impact on accuracy. Specifically, Arena achieves a maximum acceleration of end-to-end latency by 1.58×1.58$\times$1.58 × and 1.82×1.82$\times$1.82 × in two datasets, using only 54% and 34% of the bandwidth, while maintaining accuracy losses within 1% and 4% as compared to FD, respectively.
Figure 9. Visualization of Arena on two videos. In these two scenes, with a frame interval of 5, $m$ is set to 1 and 3 for MOT17 and AIC22, respectively, $p=0.9$, and $F=200$. Only the red patches in non-keyframe are used for transmission and inference.
(a) MOT17-04
(b) AIC22-S02-c006
Figure 10. Heatmaps of patches identified as PoIs, where darker areas indicate a higher frequency of offloading to the edge server.
To further illustrate the mechanism of Arena, Figure 9 presents a visualization of two scenes. The first column shows the keyframes, while the remaining four depict non-keyframes. In non-keyframes, Arena only samples an average of 22.70% and 27.79% of the total patches, yet still achieves satisfactory detection results. More visualization can be found in the supplementary materials. Additionally, Figure 10 displays the heatmaps of the PoIs extracted in these two video sequences, where darker areas indicate a higher frequency of offloading the patch to the edge server.
Moreover, Figure 11 presents the cumulative distribution function (CDF) charts of the proportion of PoI extraction in non-keyframes across various scenes. Taking the AIC22 dataset as an example, PoIs are extracted in a more granular and sparse manner, with about 80% of non-keyframes requiring the transmission of less than 25% area of the frame. This is consistent with our findings in the motivation (Sec. 2.2). The results above reveal that Arena can effectively exploit the extensive redundancy present in the video to achieve end-to-end acceleration.
Figure 11. The cumulative distribution function (CDF) of PoI proportion of all non-keyframes.
5.3. Accuracy-Bandwidth Usage Trade-off
In the design of Arena, there is a trade-off between accuracy and bandwidth usage. As depicted in Figure 12, we observe that as the keyframe interval increases, there is a reduction in bandwidth usage, but there is also a decrease in accuracy. A similar observation is made in Figure 13, where reducing the expanded bbox margin significantly decreases bandwidth usage. However, there is also a slight accuracy drop.
This trade-off relationship provides us with insights for identifying the optimal configuration. We refer to these configurations, where accuracy and bandwidth usage are well balanced, as the Sweet Spot(s). For instance, in Figure 12, these sweet spots (e.g., $K=3,4,5$) exhibit a notable reduction in bandwidth usage without a significant decrease in accuracy, thus achieving an ideal balance. Specifically, when $K=3$, Arena achieves a 38.66% reduction in bandwidth with only 0.1% loss in accuracy for MOT17Det, and a 55.76% bandwidth saving at the expense of 2.5% accuracy loss for AIC22. Additionally, as shown in Figure 13, initially (e.g., $m=0,1,2$), both accuracy and bandwidth usage increase with the expanded bbox margin $m$. However, when further increases the expanded bbox margin $m$, the accuracy does not significantly improve, thereby reaching a sweet spot. Specifically, when $m=3$, the accuracy curve of Arena reaches a stable plateau, and at this point, the bandwidth usage is 60.56% and 33.67% of the original. These optimal configurations can be obtained through offline profiling or specific strategies, which we leave for future work.
Next, we illustrate the impact of sampling rates within high probability regions on accuracy in Figure 14. For MOT17Det, accuracy remains relatively stable when the probability $p$ is within the range of 0.8 to 1.0. Similarly, for AIC22, a $p$ value exceeding 0.9 consistently keeps high accuracy. This indicates Arena’s strong capability to compensate for the input.
Figure 12. The impact of varying keyframe intervals on accuracy and bandwidth usage. ”Inf” denotes that all frames except the first one are considered as non-keyframes (i.e., $K=\infty$).
Figure 13. The impact of varying expanded bbox margin on accuracy and bandwidth usage.
Figure 14. The impact of different sampling rates on accuracy.
5.4. Model Analysis
In this subsection, we compare the model complexity of using an ViT detector with other commonly used detectors. As shown in Table 2, we benchmark the GMACs and GFlops at a resolution of $800\times 1333$ using calflops [^1] https://github.com/MrYxJ/calculate-flops.pytorch and report the frame per second (FPS). ViT-small is employed as the backbone instead of the ViT-base version in Arena because it can achieve better real-time performance while already maintaining good accuracy. When processing the videos in the dataset, Arena operates with a computing load below 25% or 50% versions in the non-keyframe inference phase. This is because the proportion of PoI in most scenes is below 50% and even 25% as shown in Figure 11.
Finally, we report the model precision in Table 3. Though Faster R-CNN offers the fastest inference speed, it requires the transmission of full frames, thus slowing down the end-to-end latency. Other models (Zhu et al., 2021; Zhang et al., 2022b) using the detection transformer (Carion et al., 2020) architecture, despite achieving commendable accuracy, fall short in real-time performance.
Table 2. Comparison of model complexity with 800 $\times$ 1333 resolution. Here we compare the inference speed (FPS) and complexity (GFlops) between widely used object detectors with different ratios of patch proportion.
Model Backbone Params (M) GMACs GFlops FPS
ViT Detector (100%) ViT-Small 41.453 222.955 446.333 21.85
ViT Detector (75%) ViT-Small 41.453 200.761 401.877 26.95
ViT Detector (50%) ViT-Small 41.453 178.589 357.463 35.26
ViT Detector (25%) ViT-Small 41.453 156.395 313.007 40.82
Faster R-CNN (Ren et al., 2015) ResNet50 41.123 214.648 430.087 47.47
Deformable-DETR (Zhu et al., 2021) ResNet50 39.823 203.641 408.499 2.73
DINO (Zhang et al., 2022b) ResNet50 47.265 286.45 574.363 2.07
  • $\triangle$ The proportion of patches used for inference.
Table 3. Model Precision
MOT17 AIC22
Model Backbone mAP@0.5 Recall mAP@0.5 Recall
ViT Detector ViT-Small 0.820 0.582 0.902 0.651
Faster R-CNN (Ren et al., 2015) ResNet50 0.791 0.587 0.865 0.620
Deformable-DETR (Zhu et al., 2021) ResNet50 0.805 0.630 0.880 0.648
DINO (Zhang et al., 2022b) ResNet50 0.819 0.632 0.876 0.656
6. Related Work
6.1. Real-Time Video Analytics Systems
Real-time video analytics is a computationally intensive task. Substantial effort focuses on enhancing the real-time performance of systems in various ways. For example, to reduce the transmission data, server-driven methods (Du et al., 2020; Zhang et al., 2022c, a) allow cameras to send low-quality videos to the edge server. The edge server then identifies RoIs and provides the camera with their position feedback. In the second transmission, only these RoIs, encoded in high quality, are sent to the server to be processed. Another line of work (Liu et al., 2019; Chen et al., 2022; Zhang et al., 2021a; Liu et al., 2022b) adopts a lightweight approach that directly extracts RoIs at the cameras and transmits only these regions to the edge server for inference. Additionally, some cloud-edge collaborative systems (Yang et al., 2023; Wang et al., 2022; Yi et al., 2020) enhance system efficiency by offloading partial computational tasks to the cloud. However, our work is the first to fully leverage characteristics of vision foundation models and achieve acceleration both on computing and transmission in the video analytics system.
6.2. Efficient Methods for Vision Transformer
Utilizing the MSA mechanism in ViT leads to computational complexity that typically scales quadratically with the number of tokens. Consequently, recent efforts have focused on reducing the number of tokens to improve the efficiency of ViT. These methods primarily fall into two categories: token pruning (Rao et al., 2021; Liang et al., 2022; Long et al., 2023; Wei et al., 2023; Yin et al., 2022; Cao et al., 2023) and token merging (Marin et al., 2023; Zeng et al., 2022; Bolya et al., 2022).
Just like what has been done to CNN-type architecture pruning in the past few years, token pruning is applied to accelerate ViT. For example, DynamicViT (Rao et al., 2021) introduces a trainable prediction module (i.e., MLP) that leverages both local and global token information to create a binary mask, which determines the token count for the next layer. Most merging-based methods rely heavily on clustering algorithms, such as K-Means (Marin et al., 2023), K-Medoids (Marin et al., 2023), and density-peak clustering with K-Nearest neighbours (Zeng et al., 2022) to cluster and merge tokens. Moreover, ToMe (Bolya et al., 2022) proposes a bipartite soft matching algorithm to cluster tokens and subsequently merge similar tokens, thus reducing the overall token count.
These ViT acceleration techniques are not suited for Arena. First, they cannot reconstruct complete feature maps, which are necessary for dense prediction tasks. Second, they require the entire frame as input. Our proposed Arena, thinking out of the box, effectively adapts efficient ViT for video analytics with its advantages.
7. Conclusion
In this paper, we find visual foundation models like ViT also have a dedicated acceleration mechanism for video analytics. Then, we design Arena, a novel edge-assisted, real-time inference acceleration system tailored for ViT-based models. This system aims to accelerate video analytics by reducing the number of patches offloaded from cameras to the server and alleviating the computation overhead of the model while still maintaining high accuracy. Our results show that Arena can accelerate the inference up to 1.58×1.58$\times$1.58 × and 1.82×1.82$\times$1.82 × on average by only using 54% and 34% bandwidth while maintaining a high inference accuracy.